# Content-based Wake-up for Energy-efficient and Timely Top-$k$ IoT Sensing Data Retrieval

### v1

@@ I Introduction
Among the possible queries, the top-$k$ query is one of the most common queries in WSNs, due to its many applications, e.g., environmental monitoring/control [9]. In most cases the data is collected to perform an action, such as controlling actuators or producing an alarm. Furthermore, such an action may often be associated with a strict time constraint, forcing the sink to obtain the top-$k$ set before a given deadline. However, the unreliable nature of the communication channel and the fact that the physical process monitored by the sensors change over time introduce a trade-off between reliability and freshness of the collected data. Specifically, if the sink requests the top-$k$ set long before the deadline, the probability that it receives all $k$ observations is large, even in the event of packet losses, etc., that necessitate retransmissions, etc. However, the physical process may change between the query instant and the deadline, resulting in a potentially large discrepancy between the measurements received by the sink and the values of the physical processes at the deadline. On the other hand, if the sink transmits the top-$k$ query close to the deadline, then there may not be sufficient time for retransmissions, etc., but the retrieved values are more likely to remain the same until the deadline. The goal of this paper is to realize both timely and energy-efficient top-$k$ data collection by exploiting wake-up radio technology.
Figure 1: An example of the top-$k$ query employing CoWu, where $k=2$: a) an example where the sink transmits CoWu signal at the early time against the deadline; b) an example where the sink transmits CoWu signal at the late time against the deadline; c) an example how the timing and the threshold of CoWu signal affects the accuracy of a top-$k$ set at the deadline.
@@ I-A Content-based Wake-up for Timely Top-$k$ Query
As argued earlier and also indicated in [8], sending a CoWu signal too early increases the risk that the collected values will be outdated at the deadline, while sending it too late may prevent some of the sensors from successfully transmitting their packets before the deadline. This is illustrated in Fig. 1(a)-(b), where a threshold $V_{\mathrm{th}}$ is transmitted using wake-up signalling $\zeta$ time slots prior to the deadline in order to collect the $k$ highest readings. In Fig. 1(a), the query is transmitted at an early time (large $\zeta$), resulting in the retrieval of all observations but with a high probability of being outdated at the deadline. On the other hand, in Fig. 1(b) the query is transmitted at a late time (small $\zeta$), resulting in only retrieving one observation, but with low probability of being outdated at the deadline.
@@ I-B1 Age of Information and Timing Requirements
AoI also has been introduced in WSNs-related research [23, 24, 25, 26, 27]. For example, in [24, 25], the authors introduced AoI in the scenario of data collection using Unmanned Aerial Vehicles (UAV), and in [26], the authors introduced AoI into sleep-wake sensors in an IIoT scenario, in which the sleep time can be set based on the physical processes and the importance of sensors with age-penalty function. In [27], authors introduced AoI in WSNs and Mobile Edge Computing (MEC).
@@ I-C Contributions and Paper Structure
  • This article introduces a new metric called $k$-QAoI for top-$k$ query with respect to a given deadline, enabling a tractable analysis of the data freshness at the deadline.
@@
  • We conduct a thorough theoretical analysis under the assumption of a realistic MAC protocol to determine the threshold of CoWu and the timing of CoWu signaling to retrieve the fresh top-$k$ set by the deadline. We further heuristically obtain the optimized set of parameters for CoWu signaling that minimizes total energy consumption while satisfying the constraint of $k$-QAoI.
@@ II-A Scenario
We study a scenario comprising a sink node and $N$ sensor nodes equipped with wake-up receivers. Each sensor, indexed by $n=1,\ldots,N$, monitors a random process $V_{n}(t)\in[V_{\mathrm{min}},V_{\mathrm{max}}]$, whose time samples are independent and identically distributed. The time is divided into recurring episodes, and each episode is divided into slots. The episodes are assumed to occur at random and infrequent instants, so that the processes monitored by the sensors can be assumed to be independent between episodes. The beginning of each episode is triggered by a request to collect data from the sensor nodes, which must be gathered before a given deadline $T$, as illustrated in Fig. 1. The requests originate from an external entity, such as an actuator or monitoring software. In this paper, we focus on top-$k$ queries, i.e., the sink aims to retrieve values of the $k$ largest readings ($k\ll N$). Given the dynamic nature of the monitored processes, collecting the top-$k$ set, i.e., the $k$ largest readings, with high freshness at the deadline is important. Note that the value of $k$ depends on the users’ requirement, whose size depends on the specific applications, and may change between episodes. For example, collecting $k$-largest values with high information freshness at the deadline is important, e.g., for controlling the $k$ number of mobile actuators or Automatic Guided Vehicles (AGV) in the factory scenario [8]. This is important in the context of natural disaster management, such as wildfire and chemical gas leakage, in which the sink needs to identify the $k$ positions of the multiple emission sources to deploy the $k$ actuators, etc., and also for multilateration techniques in the context of positioning estimation of moving objects. Furthermore, our model could be useful for some digital twin settings, where the sink needs to collect the $k$-largest readings, considering the synchronization time of the digital twin for monitoring and controlling the external entity. Here, we focus on the specific instance of a timely top-$k$ query, in which the sink knows the size of $k$ before starting the data collection.
@@ II-B1 Wake-up signal
As a wake-up control, we focus on CoWu [5], in which the threshold $V_{\mathrm{th}}$ is embedded into a wake-up signal which is transmitted to the wake-up receivers through a secondary, low-power radio link. This is achieved by encoding the estimated threshold of top-$k$ set $V_{\mathrm{th}}$ into the wake-up signal, for example, using OOK [5]. For example, the information on the threshold can be encoded into different frame lengths, in which one can map the higher (lower) threshold $V_{\mathrm{th}}$ to the shorter (longer) frame length $T_{\mathrm{wu}}$ as in [5]. This frame length can be expressed as $T_{\mathrm{wu}}(V_{\mathrm{th}})=T_{\mathrm{min}}+\mathcal{I}(V_{\mathrm{th}})T_{\mathrm{step}}$. Here, the $T_{\mathrm{min}}$ is the minimum frame, $T_{\mathrm{step}}$ is the value of the step of the different frame lengths, and $\mathcal{I}(V_{\mathrm{th}})$ is the quantization interval of threshold $V_{\mathrm{th}}$, where higher $V_{\mathrm{th}}$ corresponds to the smaller quantization interval. Note that the maximum number of bits that can be encoded into the wake-up signal is limited, such as $2^{9}$ = 512 due to the practical constraint. The possible frame length for the wake-up signal can be set by considering the standard IEEE 802.15.4g and the data size used for the data transmission [5]. Further, $T_{\mathrm{step}}$ can be chosen considering the trade-off between the resolution and the false wake-up ratio [38].
@@ II-B2 Data transmission
The sensor nodes that wake up after receiving the wake-up signal attempt data transmission through the primary radio over unlicensed frequency bands, e.g., following the IEEE 802.15.4 standard [40], and this paper assumes the use of a $p$-persistent Carrier Sense Multiple Access (CSMA) protocol for transmitting observations. Note that, as shown in [41], a $p$-persistent CSMA protocol is suited for approximating the practical MAC protocol of IEEE 802.15.4, as it exploits both the CSMA and Collision Avoidance (CA) mechanisms [5]. Therefore, each node with a relevant packet conducts carrier sensing at the beginning of a slot. Here, the duration of each slot is denoted as $\delta$ [s]. If the channel is idle, the node attempts transmission of its packet with transmission probability $p$. Each transmission attempt occupies $L$ consecutive time slots. We consider a collision channel with erasures, where collisions cause packet loss with probability one. In the absence of a collision, erasures occur with probability $e_{c}$, which is the same for all devices. An error-free Acknowledgment (ACK) is transmitted from the sink after each successful sensor data transmission. The nodes receiving an ACK transition back to sleep-state, while the nodes detecting a packet loss by the absence of an ACK retransmit their packets following $p$-persistent CSMA. Developing mechanisms to mitigate the hidden terminal problem is out of the scope of this paper and, hence, we assume that all nodes, including the sink, are located within each other’s communication, wake-up, and carrier-sensing range.
@@ II-C Energy Model
where $t_{\mathrm{tx}}^{n}$ and $t_{\mathrm{rx}}^{n}$ are time spent in transmission and receiving state for the sensor node $n$, respectively, within the episode. Note that if the sensor node $n$ remains in the sleep state during the whole episode, $t_{\mathrm{tx}}^{n}=0$ and $t_{\mathrm{rx}}^{n}=0$.
@@ II-D The $k$-QAoI Metric
To characterize the freshness of the retrieved top-$k$ set at the deadline, we introduce a new metric called $k$-QAoI, which builds upon the CoUD metric [18]. Recall that the classical AoI [13] is defined as the time elapsed since the last received packet was generated. Formally, let us denote the time slot at which $j$-th observation value of sensor node $n$ is successfully received at the sink node as $t^{\prime}_{n,j}$ and the time slot this successfully received packet is generated at the application layer as $t_{n,j}$. Then, the most recently received packet from sensor node $n$ was generated at time slot $t$
@@
To allow each episode to be analyzed in isolation, we assign a fixed age penalty $\Gamma$ to the nodes that fail to transmit their observations before the deadline. The value of $\Gamma$ corresponds to the time elapsed since the latest received packet was generated at the deadline when the nodes fail data transmission in the current episode. A small $\Gamma$ means that the sink has partial knowledge on the physical process observed the sensor nodes in the current episode based on the collected data in the previous ones, while a large $\Gamma$ means that the time between consecutive requests is sufficient and the sink has less knowledge on them.
@@ III-A CoWu
We now define the transition probability for the Markov chain. Let $P_{(m,l),(m^{{}^{\prime}},l^{{}^{\prime}})}$ be the transition probability from the state $(m,l)$ to state $(m^{{}^{\prime}},l^{{}^{\prime}})$. When the state is $(m,0)$, $m=1,2,\ldots,w$, the channel is idle and $m$ nodes attempt to transmit their packets with probability $p$. If at least one node transmits its packet, the Markov chain transitions to state $(m,1)$. Thus, this transition probability can be defined as $P_{(m,0),(m,1)}=1-(1-p)^{m}$, for $m\in\{1,2,\ldots,w\}$. On the other hand, if none of the nodes transmits, the Markov chain remains in state $(m,0)$, which happens with probability $(1-p)^{m}$. In states $(m,l)$, $l=1,\ldots,L-2$, the channel is busy, and $L-l$ slots remain of the current transmission, so the Markov chain transitions to state $(m,l+1)$ with a probability 1.
@@
$\begin{split}\ &P_{k}(r|w,w_{s})\\ &=\begin{cases}\delta(r-w_{s}),~{}~{}&\text{if}~{}w\leq k,\\ u(w-k-w_{s}+r)\frac{\binom{k}{r}\binom{w-k}{w_{s}-r}}{\binom{w}{w_{s}}},&\text{if}~{}w>k,\end{cases}\end{split}$ (12)
@@
Next, we analyze the total energy consumption of sensor nodes. We assume that the node activated by a wake-up signal continues the attempt of data transmission until it succeeds. Thus, the total energy consumption only depends on the value of $V_{\mathrm{th}}$, not on the value of $\zeta$. Here, the type of traffic considered in this paper is called One-shot Data (OSD) model [41], in which each node only has a single packet to transmit. In [41], the total energy consumption of nodes operating with $p$-persistent CSMA have been analyzed with the assumption of OSD and the collision channel, which was extended in [5], where the authors also have taken into account the packet errors due to channel impairments with an independent and identical probability of $e_{c}$ for each packet. Let us consider $w$ nodes activated by single wake-up signaling. Then, based on our energy model in Sec. II-C, $w$ out of $N$ sensor nodes consume energy, while ($N-w$) nodes remain in the sleep state. According to [5], the total energy consumption of the sensor nodes with $w$ active nodes, can be expressed as
$\begin{split}\ E_{\mathrm{total}}(w|N)&={\xi}_{T}\sum_{m=1}^{w}{\frac{L}{(1-e_{c})(1-p)^{m-1}}}{\delta}\\ &\quad+{\xi}_{R}\sum_{m=1}^{w}{\frac{L-(L-1)(1-p)^{m-1}}{(1-e_{c})p(1-p)^{m-2}}}{\delta}.\end{split}$ (14)
This work assumes $E_{\mathrm{total}}(0)=0$ as a special case. Since the number of activated nodes $w$ is a random variable, total energy consumption of each node when we employ CoWu can be expressed as follows:

### v2

@@ I Introduction
Among the possible queries, the top-$k$ query is one of the most common queries in WSNs, due to its many applications, e. g., environmental monitoring/control [9]. More generally, the sensors may compute a certain score related to the query, such as similarity, and the intention is to collect the data from the $k$ sensors that have the highest scores [10]. In most cases the data is collected to perform an action, such as controlling actuators or producing an alarm. Furthermore, such an action may often be associated with a strict time constraint, forcing the sink to obtain the top-$k$ set before a given deadline. For example, collecting the $k$ largest values with high information freshness at the deadline is important to monitor processes and identify outliers in factory scenarios or for digital twins [8]. In the context of natural disaster management, such as wildfire and chemical gas leakage detection, a timely top-$k$ query is important to identify the most critical locations to, e.g., trigger interventions. However, the unreliable nature of the communication channel and the fact that the physical process monitored by the sensors change over time introduce a trade-off between the reliability and freshness of the collected data at the deadline through the timing of sampling, data transmission, and query transmission. Specifically, if the sink requests the top-$k$ set long before the deadline, the probability that it receives all $k$ observations is large, even in the event of packet losses that necessitate retransmissions. However, the physical process may change between the query instant and the deadline, resulting in a potentially large discrepancy between the measurements received by the sink and the values of the physical processes at the deadline. On the other hand, if the sink transmits the top-$k$ query close to the deadline, then there may not be sufficient time for retransmissions but the retrieved values are more likely to remain the same until the deadline. The goal of this paper is to realize both timely and energy-efficient top-$k$ data collection by exploiting wake-up radio technology.
Figure 1: An example of the top-$k$ query employing CoWu, where $k=2$. (a): The sink transmits CoWu signal at the early time against the deadline. (b): The sink transmits CoWu signal at the late time against the deadline. (c) An example on how the timing and the threshold of CoWu signal affects the accuracy of a top-$k$ set at the deadline.
@@ I-A Content-based Wake-up for Timely Top-$k$ Query
As argued earlier and also indicated in [8], sending a CoWu signal too early increases the risk that the collected values will be outdated at the deadline, while sending it too late may prevent some of the sensors from successfully transmitting their packets before the deadline. This is illustrated in Fig. 1(a)-(b), where a threshold $V_{\mathrm{th}}$ is transmitted using wake-up signalling $\zeta$ time slots prior to the deadline in order to collect the $k$ highest readings. In Fig. 1(a), the query is transmitted at an early time (large $\zeta$), resulting in the retrieval of all observations but with a high probability of being outdated at the deadline. On the other hand, in Fig. 1(b) the query is transmitted at a late time (small $\zeta$), resulting in only retrieving one observation, but with a low probability of being outdated at the deadline.
@@ I-B1 Age of Information and Timing Requirements
AoI also has been introduced in WSNs-related research [24]. For example, in [25, 26], the authors introduced AoI in the scenario of data collection using Unmanned Aerial Vehicles (UAV), and in [27], the authors introduced AoI into sleep-wake sensors in an IIoT scenario, in which the sleep time can be set based on the physical processes and the importance of sensors with age-penalty function. In [28], authors introduced AoI in WSNs and Mobile Edge Computing (MEC).
@@ I-C Contributions and Paper Structure
  • This article introduces a new metric called $k$-QAoI for top-$k$ queries with respect to a given deadline, enabling a tractable analysis of the data freshness at the deadline.
@@
  • We conduct a thorough theoretical analysis under the assumption of a realistic MAC protocol to determine the threshold of CoWu and the timing of CoWu signaling to retrieve the fresh top-$k$ set by the deadline. We further approximately obtain the optimized set of parameters for CoWu signaling that minimizes total energy consumption while satisfying the constraint of $k$-QAoI.
@@ II-A Scenario
We study a scenario comprising a sink node and $N$ sensor nodes equipped with wake-up receivers. Each sensor, indexed by $n=1,\ldots,N$, monitors a random process $V_{n}(t)\in[V_{\mathrm{min}},V_{\mathrm{max}}]$, whose time samples are independent and identically distributed. The time is divided into recurring episodes, and each episode is divided into slots. The episodes are assumed to occur at random and infrequent instants, so that the processes monitored by the sensors can be assumed to be independent between episodes. The beginning of each episode is triggered by a request to collect data from the sensor nodes, which must be gathered before a given deadline $T$, as illustrated in Fig. 1. The requests originate from an external entity, such as an actuator or monitoring software. In this paper, we focus on top-$k$ queries, i.e., the sink aims to retrieve values of the $k$ largest readings ($k\ll N$). Given the dynamic nature of the monitored processes, collecting the top-$k$ set, i.e., the $k$ largest readings, with high freshness at the deadline is important. Note that the value of $k$ depends on the users’ requirement, whose size depends on the specific applications, and may change between episodes. Here, we focus on the specific instance of a timely top-$k$ query, in which the sink knows the size of $k$ before starting the data collection.
@@ II-B1 Wake-up signal
As a wake-up control, we focus on CoWu [5], in which the threshold $V_{\mathrm{th}}$ is embedded into a wake-up signal which is transmitted to the wake-up receivers through a secondary, low-power radio link. This is achieved by encoding the estimated threshold of the top-$k$ set $V_{\mathrm{th}}$ into the wake-up signal, for example, using OOK [5]. For example, the information on the threshold can be encoded into different frame lengths, in which one can map the higher (lower) threshold $V_{\mathrm{th}}$ to the shorter (longer) frame length $T_{\mathrm{wu}}$ as in [5]. This frame length can be expressed as $T_{\mathrm{wu}}(V_{\mathrm{th}})=T_{\mathrm{min}}+\mathcal{I}(V_{\mathrm{th}})T_{\mathrm{step}}$. Here, the $T_{\mathrm{min}}$ is the minimum frame, $T_{\mathrm{step}}$ is the value of the step of the different frame lengths, and $\mathcal{I}(V_{\mathrm{th}})$ is the quantization interval of threshold $V_{\mathrm{th}}$, where higher $V_{\mathrm{th}}$ corresponds to the smaller quantization interval. Note that the maximum number of bits that can be encoded into the wake-up signal is limited, such as $2^{9}$ = 512 due to the practical constraint. The possible frame length for the wake-up signal can be set by considering the standard IEEE 802.15.4g and the data size used for the data transmission [5]. Further, $T_{\mathrm{step}}$ can be chosen considering the trade-off between the resolution and the false wake-up ratio [39].
@@ II-B2 Data transmission
The sensor nodes that wake up after receiving the wake-up signal attempt data transmission through the primary radio over unlicensed frequency bands, e.g., following the IEEE 802.15.4 standard [41], and this paper assumes the use of a $p$-persistent Carrier Sense Multiple Access (CSMA) protocol for transmitting observations. Note that, as shown in [42], a $p$-persistent CSMA protocol is suited for approximating the practical MAC protocol of IEEE 802.15.4, as it exploits both the CSMA and Collision Avoidance (CA) mechanisms [5]. Therefore, each node with a relevant packet conducts carrier sensing at the beginning of a slot. Here, the duration of each slot is denoted as $\delta$ [s]. If the channel is idle, the node attempts transmission of its packet with transmission probability $p$. Each transmission attempt occupies $L$ consecutive time slots. We consider a collision channel with erasures, where collisions cause packet loss with probability $1$. In the absence of a collision, erasures occur with probability $e_{c}$, which is the same for all devices. An error-free Acknowledgment (ACK) is transmitted from the sink after each successful sensor data transmission. The nodes receiving an ACK transition back to sleep-state, while the nodes detecting a packet loss by the absence of an ACK retransmit their packets following $p$-persistent CSMA. Developing mechanisms to mitigate the hidden terminal problem is out of the scope of this paper and, hence, we assume that all nodes, including the sink, are located within each other’s communication, wake-up, and carrier-sensing range.
@@ II-C Energy Model
where $t_{\mathrm{tx}}^{n}$ and $t_{\mathrm{rx}}^{n}$ are time spent in transmission and receiving state for the sensor node $n$, respectively, within the episode. As described in Sec. II-B2, each wake-up node transmits data based on $p$-persistent CSMA and transitions back to the sleep state when it receives ACK. This operation leads to the different values of $t_{\mathrm{tx}}^{n}$ and $t_{\mathrm{rx}}^{n}$ for all nodes. Note that if the sensor node $n$ remains in the sleep state during the whole episode, $t_{\mathrm{tx}}^{n}=0$ and $t_{\mathrm{rx}}^{n}=0$.
@@ II-D The $k$-QAoI Metric
To characterize the freshness of the retrieved top-$k$ set at the deadline, we introduce a new metric called $k$-QAoI, which builds upon the CoUD metric [19]. Recall that the classical AoI [14] is defined as the time elapsed since the last received packet was generated. Formally, let us denote the time slot at which $j$-th observation value of sensor node $n$ is successfully received at the sink node as $t^{\prime}_{n,j}$ and the time slot this successfully received packet is generated at the application layer as $t_{n,j}$. Then, in time slot $t$, let $u_{n}(t)$ be the data generation time slot of the most recently received packet from sensor node $n$, which can be described as follows:
@@
To allow each episode to be analyzed in isolation, we assign a fixed age penalty $\Gamma$ to the nodes that fail to transmit their observations before the deadline [^4] In practice, $\Gamma$ could, for instance, be selected based on the statistics of the penalty value of all nodes, e.g., by averaging the penalty value assigned to each node based on the past data collection history. We argue that the variance of these penalties among top-$k$ nodes is small as the sink might receive data from the nodes even after it exceeds the deadline in the previous episode.. The value of $\Gamma$ corresponds to the time elapsed since the latest received packet was generated at the deadline when the nodes fail data transmission in the current episode. A small $\Gamma$ means that the sink has partial knowledge on the physical process observed the sensor nodes in the current episode based on the collected data in the previous ones, while a large $\Gamma$ means that the time between consecutive requests is sufficient and the sink has less knowledge on them.
@@
An illustration of our introduced $k$-QAoI metrics is shown in Fig. 2 for the linear age case, where the sink performs timely top-$2$ query by applying CoWu and $\Omega_{k}=\{A,B\}$. In the figure, we show the time evolution of AoI for the sensor nodes $A$ and $B$, and the time evolution of $k$-AoI, respectively. Here, the $k$-AoI can be derived by taking the average value of AoI for the nodes belonging to the top-$k$ set at the sampling time, i.e., by taking the average of AoI of both sensor nodes A and B in the example of Fig. 2. In the example of Fig. 2(a), the sink manages to collect the top-$k$ set by the deadline, which decreases AoI of sensor nodes $A$ and $B$ as well as $k$-AoI upon its successful data reception at the sink. Thanks to the successful data update by the deadline, the sink can achieve the smaller QAoI for both sensor node $A$ and sensor node $B$, leading to the smaller $k$-QAoI in Eq. (5). On the other hand, in the example of Fig. 2(b), the sink fails to collect data from sensor node $B$ by the deadline. In this case, as in the standard definition of AoI, the AoI of sensor node $B$ monotonically increases, which leads to higher QAoI. This failure of data collection by the deadline leads to higher $k$-QAoI than the case of Fig. 2(a) because the QAoI of sensor node $B$ is higher.
Figure 2: An example of the evolution of AoI, QAoI, and $k$-QAoI with linear age, where $k=2$ and $\Omega_{k}=\{A,B\}$. (a): The sink receives fresh data from nodes $A$ and $B$ by the deadline. (b): The sink receives data from sensor node $A$ while it fails to collect data from the sensor node $B$.
@@ III-A CoWu
We now define the transition probability for the Markov chain. Let $P_{(m,l),(m^{{}^{\prime}},l^{{}^{\prime}})}$ be the transition probability from the state $(m,l)$ to state $(m^{{}^{\prime}},l^{{}^{\prime}})$. When the state is $(m,0)$, $m=1,2,\ldots,w$, the channel is idle and $m$ nodes attempt to transmit their packets with probability $p$. If at least one node transmits its packet, the Markov chain transitions to state $(m,1)$. Thus, this transition probability for $m\in\{1,2,\ldots,w\}$ can be defined as
$P_{(m,0),(m,1)}=1-(1-p)^{m}.$ (9)
On the other hand, if none of the nodes transmits, the Markov chain remains in state $(m,0)$, which happens with probability $(1-p)^{m}$. In states $(m,l)$, $l=1,\ldots,L-2$, the channel is busy, and $L-l$ slots remain of the current transmission, so the Markov chain transitions to state $(m,l+1)$ with probability $1$.
@@
$\begin{split}\ &P_{k}(r|w,w_{s})\\ &=\begin{cases}\delta(r-w_{s}),~{}~{}&\text{if}~{}w\leq k,\\ u(w-k-w_{s}+r)\frac{\binom{w_{s}}{r}\binom{w-w_{s}}{k-r}}{\binom{w}{k}},&\text{if}~{}w>k,\end{cases}\end{split}$ (13)
@@
Next, we analyze the total energy consumption of sensor nodes. We assume that the node activated by a wake-up signal continues the attempt of data transmission until it succeeds even after it exceeds the deadline time following the operation of $p$-persistent CSMA. Thus, the total energy consumption only depends on the value of $V_{\mathrm{th}}$ and transmission probability $p$, but not on the value of $\zeta$. Here, the type of traffic considered in this paper is called One-shot Data (OSD) model [42], in which each node only has a single packet to transmit. In [42], the total energy consumption of nodes operating with $p$-persistent CSMA have been analyzed with the assumption of OSD and the collision channel, which was extended in [5], where the authors also have taken into account the packet errors due to channel impairments with an independent and identical probability of $e_{c}$ for each packet. Let us consider $w$ nodes activated by single wake-up signaling. Then, based on our energy model in Sec. II-C, $w$ out of $N$ sensor nodes consume energy, while ($N-w$) nodes remain in the sleep state. According to [5], the total energy consumption of the sensor nodes with $w$ active nodes, can be expressed as
$\begin{split}\ E_{\mathrm{total}}(w|N)&={\xi}_{T}\sum_{m=1}^{w}{\frac{L}{(1-e_{c})(1-p)^{m-1}}}{\delta}\\ &\quad+{\xi}_{R}\sum_{m=1}^{w}{\frac{L-(L-1)(1-p)^{m-1}}{(1-e_{c})p(1-p)^{m-2}}}{\delta},\end{split}$ (15)
where the first sum corresponds to the total energy consumed during the transmission state by $w$ wake-up nodes, and the second sum captures the total energy consumed during the receiving state by $w$ wake-up nodes, as in the energy model described in Sec. II-C. We refer the reader to [42] for a detailed derivation of Eq. (15). This work assumes $E_{\mathrm{total}}(0)=0$ as a special case. Since the number of activated nodes $w$ is a random variable, total energy consumption of each node when we employ CoWu can be expressed as follows: